marla begins walking at 3 mi/h toward the library. her friend meets her at the halfway point and drives her the rest of the way to the library. the distance to the library is 4 miles. how many hours did marla walk?
Answer by stanbon(75887)

marla begins walking at 3 mi/h toward the library. her friend meets her at the halfway point and drives her the rest of the way to the library. the distance to the library is 4 miles. how many hours did marla walk?
--------------------------
Marla walking DATA:
Rate = 3 mph ; distance = 2 miles ; time = d/r = 2/3 hour
--------------------
Marla walked for 2/3 hour
===================
Cheers,
Stan H.